OpenStudy (rsadhvika):

A small spaceship whose mass is 1.5 × 10^3 kg (including an astronaut) is drifting in outer space with negligible gravitational forces acting on it. If the astronaut turns on a 10 kW laser beam, what speed will the ship attain in 1.0 day because of the momentum carried away by the beam?

OpenStudy (samigupta8):

We can do it in this way .. The total energy carried by light beam in a day is P*t T in seconds and P in Watt.. Now this energy carried would be imparted to the ship as its final velocity So equate it with 1/2 mv^2

OpenStudy (michele_laino):

hint: the momentum transferred, is: \[\huge P = \frac{E}{c} = \frac{{W\Delta t}}{c}\] wherein \(W,\;E\) are the power and energy respectively of the laser, so the speed \(v\) of the astronaut + spaceship, is such that: \[\huge \frac{{W\Delta t}}{c} = Mv\] therefore: \[\huge v = \frac{{W\Delta t}}{{Mc}}\] where \(\Delta t=24 \times 3600= 86,400\) seconds

OpenStudy (michele_laino):

and, of course, \(c\) is the speed of light

OpenStudy (rsadhvika):

\[v = \dfrac{10000*86400}{1.5*3*10^{11}} = 1.9 mm/s\]

OpenStudy (rsadhvika):

Thank you both !

OpenStudy (michele_laino):

:)

OpenStudy (samigupta8):

@rsadhvika i were wrong for i didn't see the last line of your question... The momentum one...

OpenStudy (samigupta8):

And i really don't deserve the medal...so pls..give it to michele sir